How much water must be evaporated from 100 gallongs of a 4% brine solution to get a 5% brine solution?
Answer by Alan3354(69443)

How much water must be evaporated from 100 gallongs of a 4% brine solution to get a 5% brine solution?
-----------------
4% has 4 gallons of salt.
For 4 gallons to be 5%, Total solution = 4/0.05 = 80 gallons
100 - 80 = 20 gallons to be evaporated.